# Nearly Minimax Optimal Regret for Multinomial Logistic Bandit

### v6

@@ 1 Introduction
The multinomial logistic (MNL) bandit framework [46, 47, 6, 7, 39, 40, 43, 4, 52] describes sequential assortment selection problems in which an agent offer a sequence of assortments of at most $K$ item from a set of $N$ possible items and receives feedback only for the chosen decisions. The choice probability of each outcome is characterized by an MNL model [36]. This framework allows modeling of various real-world situations such as recommender systems and online retails, where selections of assortments are evaluated based on the user-choice feedback among offered multiple options.
In this paper, we study the contextual MNL bandit problem [7, 6, 42, 14, 39, 40, 43, 4], where the features of items and possibly contextual information about a user at each round are available. Despite many recent advances, [14, 39, 40, 43, 4], however, no previous studies have proven the minimax optimality of contextual MNL bandits. Chen et al. [14] proposed a regret lower bound of $\Omega(d\sqrt{T}/K)$, where $d$ is the number of features, $T$ is the total number of rounds, and $K$ is the maximum size of assortments, assuming the uniform rewards, i.e., rewards are all same for each of the total $N$ items. Furthermore, Chen and Wang [13] established a regret lower bound of $\Omega(\sqrt{NT})$ in the non-contextual setting (hence, dependence on $N$ appears instead of $d$), which is tighter in terms of $K$. It is important to note the difference in the assumptions for the attraction parameter for the outside option $v_{0}$. Chen and Wang [13] assumed for the attraction parameter for the outside option to be $v_{0}=K$, whereas Chen et al. [14] assumed $v_{0}=1$. Therefore, it remains an open question whether and how the value of $v_{0}$ affects both lower and upper bounds of regret.
@@
Table 1: Comparisons of lower and upper regret bounds in related works on MNL bandits with $T$ rounds, $N$ items, the maximum size of assortments $K$, $d$-dimensional feature vectors, and problem-dependent constants $1/\kappa=\mathcal{O}(K^{2})$ and $\kappa^{\prime}=\mathcal{O}(1/K)$. $\tilde{\mathcal{O}}$ represents big-$\mathcal{O}$ notation up to logarithmic factors. For the computational cost (abbreviated as “Comput.”), we consider only the dependence on the number of rounds $t$. “Intractable” means a non-polynomial runtime. The notation “--” denotes not applicable. The starred () papers only consider the non-contextual setting.
@@
Lower Bound Chen et al. [14] $\Omega(d\sqrt{T}/K)$- Uniform $\Theta(1)$-
@@
This work (Theorem 1) $\Omega(\frac{\sqrt{v_{0}K}}{v_{0}+K}d\sqrt{T})$- Uniform Any value --
@@
Upper Bound Chen et al. [14] $\tilde{\mathcal{O}}(d\sqrt{T})$ Stochastic Non-uniform $\Theta(1)$ Intractable
@@ 3 Problem Setting
Notations. For a positive integer, $n$, we denote $[n]:=\{1,2,\ldots,n\}$. For a real-valued matrix $A$, we denote $\|A\|_{2}:=\sup_{x:\|x\|_{2}=1}\|Ax\|_{2}$ as the maximum singular value of $A$. For two symmetric matrices, $V$ and $W$ of the same dimensions, $V\succeq W$ means that $V-W$ is positive semi-definite. Finally, we define $\mathcal{S}$ to be the set of candidate assortment with size constraint at most $K$, i.e., $\mathcal{S}=\{S\subset[N]:|S|\leq K\}$. While, for simplicity, we consider both $\mathcal{S}$ and the set of items $[N]$ to be stationary in this paper, it is important to note that both $\mathcal{S}$ and $[N]$ can vary over time.
Contextual MNL bandits. We consider a sequential assortment selection problem which is defined as follows. At each round $t$, the agent observes feature vectors $x_{ti}\in\mathbb{R}^{d}$ for every item $i\in[N]$. Based on this contextual information, the agent presents an assortment $S_{t}=\{i_{1},\dots,i_{l}\}\in\mathcal{S}$, where $l\leq K$, and then observes the user purchase decision $c_{t}\in S_{t}\cup\{0\}$, where $\{0\}$ represents the “outside option” which indicates that the user did not select any of the items in $S_{t}$. The distribution of these selections follows a multinomial logit (MNL) choice model [36], where the probability of choosing any item $i_{k}\in S_{t}$ (or the outside option) is defined as:
@@
In the existing literature on MNL bandits, it is commonly assumed that $v_{0}=1$ [39, 40, 43, 4, 52]. On the other hand, Chen and Wang [13], Agrawal et al. [7] assume that $v_{0}=K$ 111 Chen and Wang [13] indeed set $v_{0}=1$ and $v_{1},\dots,v_{N}=\Theta(1/K)$. However, this is equivalent to the setting with $v_{0}=K$ and $v_{1},\dots,v_{N}=\Theta(1)$. to induce a tighter lower bound in terms of $K$. Later, we will explore how these differing assumptions create fundamentally different problems, leading to different regret lower bounds (Subsection 5.1).
@@
There exist $\kappa>0$ such that for every item $i\in S$ and any $S\in\mathcal{S}$, and all round $t$, $\min_{\mathbf{w}\in\mathcal{W}}p_{t}(i|S,\mathbf{w})p_{t}(0|S,\mathbf{w})\geq\kappa$, where $\mathcal{W}=\{\mathbf{w}\in\mathbb{R}^{d}\mid\|\mathbf{w}\|_{2}\leq 1\}$.
@@ 5.1 Regret Lower Bound under Uniform Rewards
Let $d$ be divisible by $4$ and let Assumption 1 hold true. Suppose $T\geq C\cdot d^{4}(v_{0}+K)/K$ for some constant $C>0$. Then, in the uniform reward setting, for any policy $\pi$, there exists a worst-case problem instance with $N=\Theta(K\cdot 2^{d})$ items such that the worst-case expected regret of $\pi$ is lower bounded as follows:
@@ 5.2 Minimax Optimal Regret Upper Bound under Uniform Rewards
Online parameter estimation. Instead of performing MLE as in previous works [14, 40, 43], inspired by Zhang and Sugiyama [52], we use the online mirror descent algorithm to estimate the parameter. We first define the multinomial logistic loss function at round $t$ as:
@@
and $H_{t}:=\lambda\mathbf{I}_{d}+\sum_{s=1}^{t-1}\mathcal{G}_{s}(\mathbf{w}_{s+1})$. Note that $\mathcal{G}_{t}(\mathbf{w})=\nabla^{2}\ell_{t}(\mathbf{w})$. This online estimator is efficient in terms of both computation and storage. By a standard online mirror descent formulation [41], the optimization problem in (3) can be solved using a single projected gradient step through the following equivalent formula:
@@
2:Initialize: $H_{1}=\lambda\mathbf{I}_{d}$ and $\mathbf{w}_{1}$ as any point in $\mathcal{W}$,
@@
6:     Update $\tilde{H}_{t}=H_{t}+\eta\mathcal{G}_{t}(\mathbf{w}_{t})$, and update the estimator $\mathbf{w}_{t+1}$ by (3).
@@
Discussion of Theorem 2. If $T\geq\mathcal{O}(d^{2}(v_{0}+K)^{2}/(\kappa^{2}v_{0}K))$, Theorem 2 shows that our algorithm OFU-MNL+ achieves minimax optimal regret (up to logarithmic factor) in terms of all $d$, $T$, $K$, and even $v_{0}$. To the best of our knowledge, ignoring logarithmic factors, our proposed algorithm is the first computationally efficient, minimax optimal algorithm in (adversarial) contextual MNL bandits. When $v_{0}=\Theta(1)$, which is the convention in existing MNL bandit literature [39, 40, 43, 4, 52]OFU-MNL+ obtains $\tilde{\mathcal{O}}(d\sqrt{T/K})$ regret. This represents an improvement over the previous upper bound of Perivier and Goyal [43] 222 Perivier and Goyal [43] also consider the uniform rewards ($r_{ti}=1$) with $v_{0}=1$., which is $\tilde{\mathcal{O}}(dK\sqrt{\kappa^{\prime}T}+d^{2}K^{4}/\kappa)$, where $\kappa^{\prime}=\mathcal{O}(1/K)$, by a factor of $K$. This improvement can be attributed to two key factors: an improved, constant, self-concordant-like property of the loss function (Proposition C.1) and a $K$-free elliptical potential lemma (Lemma E.2). Furthermore, by employing an improved bound for the second derivative of the revenue (Lemma E.3), we achieve an enhancement in the regret for the second term, $d^{2}/\kappa$, by a factor of $K^{4}$, in comparison to Perivier and Goyal [43]. Unless $v_{0}=\Theta(K)$, Theorem 2 indicates that the regret decreases as the assortment size $K$ increases. To the best of our knowledge, this is the first algorithm in MNL bandits to show that increasing $K$ results in a reduction in regret. Moreover, when reduced to the logistic bandit, i.e., $K=1$, $r_{t1}=1$, and $v_{0}=1$, our algorithm can also achieve a regret of $\tilde{\mathcal{O}}(d\sqrt{\kappa T})$ by Corollary 1 in Zhang and Sugiyama [52], which is consistent with the results in Abeille et al. [3], Faury et al. [21]. The proof is deferred to Appendix E.
@@
The proposed algorithm is computationally efficient in both parameter updates and assortment selections. Since we employ online parameter estimation, akin to Zhang and Sugiyama [52], our algorithm demonstrates computational efficiency in parameter estimation, incurring only incurring $\mathcal{O}(Kd^{3})$ computation cost and $\mathcal{O}(d^{2})$ storage cost, which are completely independent of $t$. Furthermore, a naive approach to selecting the optimistic assortment requires enumerating all possible ($N$ choose $K$) assortments, resulting in exponential computational cost [14]. However, by constructing the optimistic expected revenue according to (6) (inspired by Oh and Iyengar [40]), our algorithm needs only $\mathcal{O}(N)$ computational cost.
@@ 5.3 Regret Upper & Lower Bounds under Non-Uniform Rewards
Under the same conditions as Theorem 1, let the rewards be non-uniform and $v_{0}=\Theta(1)$. Then, for any policy $\pi$, there exists a worst-case problem instance such that the worst-case expected regret of $\pi$ is lower bounded as follows:
@@
We also prove a matching upper bound up to logarithmic factors. The algorithm OFU-MNL+ is also applicable in the case of non-uniform rewards. However, because the optimistic expected revenue $\tilde{R}_{t}(S)$ is no longer an increasing function of $\alpha_{ti}$, optimizing for $S_{t}=\mathop{\mathrm{argmax}}_{S\in\mathcal{S}}\tilde{R}_{t}(S)$ no longer equates to simply selecting the top $K$ items with the highest optimistic utility. Instead, we employ assortment optimization methods introduced in Rusmevichientong et al. [46], Davis et al. [19], which are efficient polynomial-time (independent of $t$ [^3] An interior point method would generally solve the problem with a computational complexity of $\mathcal{O}(N^{3.5})$. algorithms available for solving this optimization problem. Therefore, our algorithm is also computationally efficient under non-uniform rewards.
@@ 7 Numerical Experiments
In this section, we empirically evaluate the performance of our algorithm OFU-MNL+. We measure cumulative regret over $T=3000$ rounds. For each experimental setup, we run the algorithms across $20$ independent instances and report the average performance. In each instance, the underlying parameter $\mathbf{w}^{\star}$ is randomly sampled from a $d$-dimensional uniform distribution, where each element of $\mathbf{w}^{\star}$ lies within the range $[-1/\sqrt{d},1/\sqrt{d}]$ and is not known to the algorithms. Additionally, the context features $x_{ti}$ are drawn from a $d$-dimensional multivariate Gaussian distribution, with each element of $x_{ti}$ clipped to the range $[-1/\sqrt{d},1/\sqrt{d}]$. This setup ensures compliance with Assumption 1. In the uniform reward setting (first row of Figure 1), the combinatorial optimization step to choose the assortment reduces to sorting items by their utility estimate. In the non-uniform reward setting (second row of Figure 1), the rewards are sampled from a uniform distribution in each round, i.e., $r_{ti}\sim\operatorname{Unif}(0,1)$. Refer Appendix K for more details.
@@ Appendix B Notation
$R_{t}(S,\mathbf{w}^{\star})$ $:=\sum_{i\in S}p_{t}(i|S,\mathbf{w}^{\star})r_{ti}$, expected revenue of the assortment $S$ at round $t$
$\ell_{t}(\mathbf{w})$ $:=-\sum_{i\in S_{t}}y_{ti}\log\left(\frac{\exp(x_{ti}^{\top}\mathbf{w})}{v_{0}\!+\!\sum_{j\in S_{t}}\exp(x_{tj}^{\top}\mathbf{w})}\right)$, loss function at round $t$
$\ell(\mathbf{z}_{t},\mathbf{y}_{t})$ $:=-\sum_{i\in S_{t}}y_{ti}\log\left(\frac{\exp(z_{ti})}{v_{0}\!+\!\sum_{j\in S_{t}}\exp(z_{tj})}\right)$, loss function at round $t$, $z_{ti}=x_{ti}^{\top}\mathbf{w}$
@@
$\mathcal{G}_{t}(\mathbf{w})$ $:=\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w})x_{ti}x_{ti}^{\top}-\sum_{i\in S_{t}}\sum_{j\in S_{t}}p_{t}(i|S_{t},\mathbf{w})p_{t}(j|S_{t},\mathbf{w})x_{ti}x_{tj}^{\top}$
$H_{t}$ $:=\lambda\mathbf{I}_{d}+\sum_{s=1}^{t-1}\mathcal{G}_{s}(\mathbf{w}_{s+1})$
$\tilde{H}_{t}$ $:=H_{t}+\eta\mathcal{G}_{t}(\mathbf{w}_{t})$
$\alpha_{ti}$ $:=x_{ti}^{\top}\mathbf{w}_{t}+\beta_{t}(\delta)\|x_{ti}\|_{H_{t}^{-1}}$, optimistic utility for item $i$ at round $t$
$\beta_{t}(\delta)$ $:=\mathcal{O}\left(\sqrt{d}\log t\log K\right)$, confidence radius at round $t$
$\tilde{R}_{t}(S)$ $:=\frac{\sum_{i\in S}\exp(\alpha_{ti})r_{ti}}{v_{0}+\sum_{j\in S}\exp(\alpha_{tj})}$, optimistic expected revenue for the assortment $S$ at round $t$

### v9

@@ 1 Introduction
The multinomial logistic (MNL) bandit framework [45, 46, 6, 7, 38, 39, 42, 4, 51] describes sequential assortment selection problems in which an agent offer a sequence of assortments of at most $K$ items from a set of $N$ possible items and receives feedback only for the chosen decisions. The choice probability of each outcome is characterized by an MNL model [35]. This framework allows modeling of various real-world situations such as recommender systems and online retails, where selections of assortments are evaluated based on the user-choice feedback among offered multiple options.
In this paper, we study the contextual MNL bandit problem [7, 6, 41, 14, 38, 39, 42, 4], where the features of items and possibly contextual information about a user at each round are available. Despite many recent advances, [14, 38, 39, 42, 4], however, no previous studies have proven the minimax optimality of contextual MNL bandits. Chen et al. [14] proposed a regret lower bound of $\Omega(d\sqrt{T}/K)$, where $d$ is the feature dimension, $T$ is the total number of rounds, and $K$ is the maximum size of assortments, assuming the uniform rewards, i.e., rewards are all same for each of the total $N$ items. Furthermore, Chen and Wang [13] established a regret lower bound of $\Omega(\sqrt{NT})$ in the non-contextual setting (hence, dependence on $N$ appears instead of $d$), which is tighter in terms of $K$. It is important to note the difference in the assumptions for the attraction parameter for the outside option $v_{0}$. Chen and Wang [13] assumed for the attraction parameter for the outside option to be $v_{0}=K$, whereas Chen et al. [14] assumed $v_{0}=1$. Therefore, it remains an open question whether and how the value of $v_{0}$ affects both lower and upper bounds of regret.
@@
Table 1: Comparisons of lower and upper regret bounds in related works on MNL bandits with $T$ rounds, NN items, the maximum size of assortments $K$, $d$-dimensional feature vectors, and problem-dependent constants $1/\kappa=\mathcal{O}(K^{2})$ and $\kappa^{\prime}=\mathcal{O}(1/K)$. $\tilde{\mathcal{O}}$ represents big-$\mathcal{O}$ notation up to logarithmic factors. For the computational cost (abbreviated as “Comput.”), we consider only the dependence on the number of rounds $t$. “Intractable” means a non-polynomial runtime. The notation “-” denotes not applicable. The starred () papers only consider the non-contextual setting.
@@
Lower
Bound
Chen et al. [14] $\Omega(d\sqrt{T}/K)$ - Uniform $\Theta(1)$ -
@@
This work (Theorem 1) $\Omega(\frac{\sqrt{v_{0}K}}{v_{0}+K}d\sqrt{T})$ - Uniform Any value -
@@
Upper
Bound
Chen et al. [14] $\tilde{\mathcal{O}}(d\sqrt{T})$ Stochastic Non-uniform $\Theta(1)$ Intractable
@@ 3 Problem Setting
Notations. For a positive integer, $n$, we denote $n]:=\{1,2,\ldots,n\}$. For a real-valued matrix $A$, we denote $\|A\|_{2}:=\sup_{x:\|x\|_{2}=1}\|Ax\|_{2}$ as the maximum singular value of $A$. For two symmetric matrices, $V$ and $W$ of the same dimensions, $V\succeq W$ means that VWV-W is positive semi-definite. Finally, we define $\mathcal{S}$ to be the set of candidate assortment with size constraint at most $K$, i.e., $\mathcal{S}=\{S\subseteq[N]:|S|\leq K\}$. While, for simplicity, we consider both $\mathcal{S}$ and the set of items $[N]$ to be stationary in this paper, it is important to note that both $\mathcal{S}$ and $[N]$ can vary over time.
Contextual MNL bandits. We consider a sequential assortment selection problem which is defined as follows. At each round $t$, the agent observes feature vectors $x_{ti}\in\mathbb{R}^{d}$ for every item $i\in[N]$. Based on this contextual information, the agent presents an assortment $S_{t}=\{i_{1},\dots,i_{l}\}\in\mathcal{S}$, where $l\leq K$, and then observes the user purchase decision $c_{t}\in S_{t}\cup\{0\}$, where $\{0\}$ represents the “outside option” which indicates that the user did not select any of the items in $S_{t}$. The distribution of these selections follows a multinomial logistic (MNL) choice model [35], where the probability of choosing any item $i_{k}\in S_{t}$ (or the outside option) is defined as:
@@
In the existing literature on MNL bandits, it is commonly assumed that $v_{0}=1$ [38, 39, 42, 4, 51]. On the other hand, Chen and Wang [13], Agrawal et al. [7] assume that $v_{0}=K$  [^1] Chen and Wang [13] indeed set $v_{0}=1$ and $v_{1},\dots,v_{N}=\Theta(1/K)$. However, this is equivalent to the setting with $v_{0}=K$ and $v_{1},\dots,v_{N}=\Theta(1)$. to induce a tighter lower bound in terms of $K$. Later, we will explore how these differing assumptions create fundamentally different problems, leading to different regret lower bounds (Subsection 5.1).
@@
There exists $\kappa>0$ such that for every item $i\in S$, any $S\in\mathcal{S}$, and all round $t$, $\min_{\mathbf{w}\in\mathcal{W}}p_{t}(i|S,\mathbf{w})p_{t}(0|S,\mathbf{w})\geq\kappa$, where $\mathcal{W}=\{\mathbf{w}\in\mathbb{R}^{d}\mid\|\mathbf{w}\|_{2}\leq 1\}$.
@@ 5.1 Regret Lower Bound under Uniform Rewards
Let $d$ be divisible by $4$ and let Assumption 1 hold true. Suppose $T\geq C\cdot d^{4}(v_{0}+K)/K$ for some constant $C>0$. Then, in the uniform reward setting, for any policy $\pi$, there exists a worst-case problem instance such that the worst-case expected regret of $\pi$ is lower bounded as follows:
@@ 5.2 Minimax Optimal Regret Upper Bound under Uniform Rewards
Online parameter estimation. Instead of performing MLE as in previous works [14, 39, 42], inspired by Zhang and Sugiyama [51], we use the online mirror descent algorithm to estimate parameter. We first define the multinomial logistic loss function at round $t$ as:
@@
and $H_{t}:=\lambda\mathbf{I}_{d}+\sum_{s=1}^{t-1}\mathcal{G}_{s}(\mathbf{w}_{s+1})$. Note that $\mathcal{G}_{t}(\mathbf{w})=\nabla^{2}\ell_{t}(\mathbf{w})$. This online estimator is efficient in terms of both computation and storage. By a standard online mirror descent formulation [40], the optimization problem in Equation (3) can be solved using a single projected gradient step through the following equivalent formula:
@@
2:Initialize: $H_{1}=\lambda\mathbf{I}_{d}$ and $\mathbf{w}_{1}$ as any point in $\mathcal{W}$.
@@
6:  Update $\tilde{H}_{t}=H_{t}+\eta\mathcal{G}_{t}(\mathbf{w}_{t})$, and update the estimator $\mathbf{w}_{t+1}$ by Equation (3).
@@
Discussion of Theorem 2. If $T\geq\mathcal{O}(d^{2}(v_{0}+K)^{2}/(\kappa^{2}v_{0}K))$, Theorem 2 shows that our algorithm OFU-MNL+ achieves minimax optimal regret (up to logarithmic factor) in terms of all $d$, $T$, $K$, and even $v_{0}$. To the best of our knowledge, ignoring logarithmic factors, our proposed algorithm is the first computationally efficient, minimax optimal algorithm in (adversarial) contextual MNL bandits. When $v_{0}=\Theta(1)$, which is the convention in existing MNL bandit literature [38, 39, 42, 4, 51], OFU-MNL+ obtains $\tilde{\mathcal{O}}(d\sqrt{T/K})$ regret. This represents an improvement over the previous upper bound of Perivier and Goyal [42] 222 Perivier and Goyal [42] also consider the uniform rewards ($r_{ti}=1$) with $v_{0}=1$., which is $\tilde{\mathcal{O}}(dK\sqrt{\kappa^{\prime}T}+d^{2}K^{4}/\kappa)$, where $\kappa^{\prime}=\mathcal{O}(1/K)$, by a factor of $K$. This improvement can be attributed to two key factors: an improved, constant, self-concordant-like property of the loss function (Proposition C.1) and a $K$-free elliptical potential lemma (Lemma E.2). Furthermore, by employing an improved bound for the second derivative of the revenue (Lemma E.3), we achieve an enhancement in the regret for the second term, $d^{2}/\kappa$, by a factor of $K^{4}$, in comparison to Perivier and Goyal [42]. Unless $v_{0}=\Theta(K)$, Theorem 2 indicates that the regret decreases as the assortment size $K$ increases. To the best of our knowledge, this is the first algorithm in MNL bandits to show that increasing $K$ results in a reduction in regret. Moreover, when reduced to the logistic bandit, i.e., $K=1$, $r_{t1}=1$, and $v_{0}=1$, our algorithm can also achieve a regret of $\tilde{\mathcal{O}}(d\sqrt{\kappa T})$ by Corollary 1 in Zhang and Sugiyama [51], which is consistent with the results in Abeille et al. [3], Faury et al. [21]. The proof is deferred to Appendix E.
@@
The proposed algorithm is computationally efficient in both parameter updates and assortment selections. Since we employ online parameter estimation, akin to Zhang and Sugiyama [51], our algorithm demonstrates computational efficiency in parameter estimation, incurring only incurring $\mathcal{O}(Kd^{3})$ computation cost and $\mathcal{O}(d^{2})$ storage cost, which are completely independent of $t$. Furthermore, a naive approach to selecting the optimistic assortment requires enumerating all possible ($N$ choose $K$) assortments, resulting in exponential computational cost [14]. However, by constructing the optimistic expected revenue according to Equation (6) (inspired by Oh and Iyengar [39]), our algorithm needs only $\mathcal{O}(N)$ computational cost.
@@ 5.3 Regret Upper & Lower Bounds under Non-Uniform Rewards
Suppose Assumption 1 holds and $T\geq C\cdot d^{4}(v_{0}+1)/1$ for some constant $C>0$. Let $d$ be divisible by $4$ and $v_{0}=\Theta(1)$. Then, under the non-uniform reward setting, for any policy $\pi$, there exists an instance such that the worst-case expected regret of $\pi$ is lower bounded as follows:
@@
We also prove a matching upper bound up to logarithmic factors. The algorithm OFU-MNL+ is also applicable in the case of non-uniform rewards. However, because the optimistic expected revenue $\tilde{R}_{t}(S)$ is no longer an increasing function of $\alpha_{ti}$, optimizing for $S_{t}=\mathop{\mathrm{argmax}}_{S\in\mathcal{S}}\tilde{R}_{t}(S)$ no longer equates to simply selecting the top $K$ items with the highest optimistic utility. Instead, we employ assortment optimization methods introduced in Rusmevichientong et al. [45], Davis et al. [19], which are efficient polynomial-time (independent of $t$)  [^3] An interior point method would generally solve the problem with a computational complexity of $\mathcal{O}(N^{3.5})$. algorithms available for solving this optimization problem. Therefore, our algorithm is also computationally efficient under non-uniform rewards.
@@ 7 Numerical Experiments
In this section, we empirically evaluate the performance of our algorithm OFU-MNL+. We measure cumulative regret over $T=3000$ rounds. For each experimental setup, we run the algorithms across $20$ independent instances and report the average performance. In each instance, the underlying parameter $\mathbf{w}^{\star}$ is randomly sampled from a $d$-dimensional uniform distribution, where each element of $\mathbf{w}^{\star}$ lies within the range $1/\sqrt{d},1/\sqrt{d}]$ and is not known to the algorithms. Additionally, the context features $x_{ti}$ are drawn from a $d$-dimensional multivariate Gaussian distribution, with each element of $x_{ti}$ clipped to the range $1/\sqrt{d},1/\sqrt{d}]$. This setup ensures compliance with Assumption 1. In the uniform reward setting (first row of Figure 1), the combinatorial optimization step to choose the assortment reduces to sorting items by their utility estimate. In the non-uniform reward setting (second row of Figure 1), the rewards are sampled from a uniform distribution in each round, i.e., $r_{ti}\sim\operatorname{Unif}(0,1)$. Refer Appendix K for more details.
@@ Appendix B Notation
$R_{t}(S,\mathbf{w}^{\star})$ $\sum_{i\in S}p_{t}(i|S,\mathbf{w}^{\star})r_{ti}$, expected revenue of the assortment $S$ at round $t$
$\ell_{t}(\mathbf{w})$ $\sum_{i\in S_{t}}y_{ti}\log\left(\frac{\exp(x_{ti}^{\top}\mathbf{w})}{v_{0}\!+\!\sum_{j\in S_{t}}\exp(x_{tj}^{\top}\mathbf{w})}\right)$, loss function at round $t$
$\ell(\mathbf{z}_{t},\mathbf{y}_{t})$ $\sum_{i\in S_{t}}y_{ti}\log\left(\frac{\exp(z_{ti})}{v_{0}\!+\!\sum_{j\in S_{t}}\exp(z_{tj})}\right)$, loss function at round $t$, $z_{ti}=x_{ti}^{\top}\mathbf{w}$
@@
$\mathcal{G}_{t}(\mathbf{w})$ $\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w})x_{ti}x_{ti}^{\top}-\sum_{i\in S_{t}}\sum_{j\in S_{t}}p_{t}(i|S_{t},\mathbf{w})p_{t}(j|S_{t},\mathbf{w})x_{ti}x_{tj}^{\top}$
$H_{t}$ $\lambda\mathbf{I}_{d}+\sum_{s=1}^{t-1}\mathcal{G}_{s}(\mathbf{w}_{s+1})$
$\tilde{H}_{t}$ $H_{t}+\eta\mathcal{G}_{t}(\mathbf{w}_{t})$
$\alpha_{ti}$ $x_{ti}^{\top}\mathbf{w}_{t}+\beta_{t}(\delta)\|x_{ti}\|_{H_{t}^{-1}}$, optimistic utility for item $i$ at round $t$
$\beta_{t}(\delta)$ $\mathcal{O}\left(\sqrt{d}\log t\log K\right)$, confidence radius at round $t$
$\tilde{R}_{t}(S)$ $\frac{\sum_{i\in S}\exp(\alpha_{ti})r_{ti}}{v_{0}+\sum_{j\in S}\exp(\alpha_{tj})}$, optimistic expected revenue for the assortment $S$ at round $t$